OpenStudy (anonymous):

Please helppp: You buy a car for $13,000 and take out a loan for the whole thing with no down payment. If the loan is paid over five years at 9%, what is the monthly payment?

jimthompson5910 (jim_thompson5910):

you will use the formula P = L[c(1 + c)^n]/[(1 + c)^n - 1] where P = monthly payment L = total loan amount c = monthly interest rate n = number of payments

jimthompson5910 (jim_thompson5910):

In this case, P = unknown L = 13000 c = 0.09/12 = 0.0075 n = 12*5 = 60